Ace 112 Project 10 Following are independent situations:

Ace 112 Project 10

Following are independent situations:


Nicholas Ram Corporation have a $2,400,000 "bond issue" dated March 1, 2016 due in 15 years with an annual interest rate of 10%. Interest is payable March 1 and September 1. On August 1, 2016, the bond was sold for $2,478,750 plus accrued interest. 

Using the straight-line method, prepare the general journal entries for each of the following 

a) The issuance of the bond on August 1, 2016 

b) Payment of the semi-annual interest and the amortization of the premium on September 1, 2016 

c) Accrual of the interest and the amortization of the premium on December 31, 2016 

d) Payment of the semi-annual interest and the amortization of the premium on March 1, 2017
Answer #1

Dear student, we cannot able to post solution more than one question as per our policy.

Clean price of bonds issued $            2,478,750
Less: face value of bonds $            2,400,000
Premium on bonds payable $                  78,750
Clean price of bonds issued $            2,478,750
Plus: Accured interest (March 1 to August 1 = 5 months) (2400000*10%*5/12) $               100,000
Total cash received $            2,578,750
Premium on bonds payable $                  78,750
Divided by: Number of months (15*12 = 180) (180-5 months already gone. = 175)                            175
Amortization of premium per month $                        450
Journal entries
Date Account title Debit Credit
2016
Aug 1 Cash $      2,578,750
Premium on bonds payable $              78,750
Bonds payable $        2,400,000
Bonds interest payable $            100,000
To record the issue of bonds
Sep 1 Bonds interest expense $            19,550
Bonds interest payable $          100,000
Premium on bonds payable (450*1 month) $                  450
Cash (2400000*10%*6/12) $            120,000
To record the interest paid on bonds
Dec 31 Bonds interest expense $            78,200
Premium on bonds payable (450*4 months) $               1,800
Bonds interest payable (2400000*10%*4/12) $              80,000
To record the accrued interest on bonds ( Sep 1 to Dec 31 = 4 months)
2017
Mar 1 Bonds interest expense $            39,100
Bonds interest payable $            80,000
Premium on bonds payable (450*2 month) $                  900
Cash (2400000*10%*6/12) $            120,000
To record the interest paid on bonds (Jan 1 to Mar 1 = 2 months for interest expense)